OpenStudy (ashking1):

What is the simplified form of the expression 8 sqrt v^2-sqrt v^4-sqrt 9v^2+sqrt 16v4

OpenStudy (anonymous):

\[\large 8\sqrt{v^2}-\sqrt{v^4}-\sqrt{9v^2}+\sqrt{16v^4} \] this?

OpenStudy (ashking1):

yes

OpenStudy (anonymous):

\[\large 8\sqrt{v^2}-\sqrt{v^4}-\sqrt{9v^2}+\sqrt{16v^4}=8v-v^2-3v+4v^2 \] can you take it from here?

OpenStudy (ashking1):

no im still stuck

OpenStudy (anonymous):

combine the v^2 terms: 4v^2 - v^2 combine the v terms: 8v - 3v string the results with a "+"

OpenStudy (ashking1):

4v+5v

OpenStudy (anonymous):

no... try again...

OpenStudy (ashking1):

3v^2+5v